TiMoThYnICee:

Hello everyone I'm new here and I already know this answer I just want everyone else to know as well. What is 12.5% as a fraction

TiMoThYnICee:

The answer would be 1/8. 12.5/100 = 0.125 0.125/1 x 1000/1000 = 125/1000 and simplify that and you get 1/8.

TiMoThYnICee:

that's the answer hope I helped